# JoPA: Explaining Large Language Model’s Generation via Joint Prompt Attribution

### v2

@@ A.4 Benefits to the community on improving model ability
JoPA could be used as the defense of the jailbreak attacks and help users to design better prompts to get satisfying responses. We also do the experiments to show the effectiveness of applying JoPA to detect malicious prompts on Llama-2 (7B-Chat) model, which is similar to the application done in CONTEXTCITE (contextcite, ) The results are shown below:
@@ A.6 Ablation Study
In addition, we extend the application of our framework, JoPA, to a larger model and a more challenging task, specifically Few-shot Chain-of-Thought (CoT) reasoning (wei2022chain, ; kojima2022large, ), as described in Appendix A.11 and Appendix A.12. The better performance of our algorithm compared with others highlights the transferability of our framework across different model architectures and tasks.
@@ A.12 Method Generalizability on Reasoning Task
We demonstrate the generalizability of our algorithm by conducting an additional experiment on the reasoning task of Few-shot-CoT (wei2022chain, ; kojima2022large, ), using the dataset AQuA (kojima2022large, ; goswami2024aqua, ). The following results in Figure 8 shows the remarkable performance of our method, indicating its ability to generalize well on more complex tasks like CoT.
@@ A.13 Comparison with Other Baselines
ReAGent (zhao2024reagent, ) addresses a different task than JoPA. It focuses on classification tasks by explaining the importance of words in inputs corresponding to the predicted label and extends this approach to generation tasks by explaining the single next predicted word. In contrast, our task focuses on explaining the relationship between the input prompt and the entire generated sentences. To compare the performance of JoPA and ReAGent on the generation task, we modify ReAGent’s explanatory target to encompass the full generation outputs rather than a single word. We conduct experiments on the Alpaca dataset using the Llama-2 (7B-Chat) model, and the results are presented in Figure 9. The experimental results demonstrate that JoPA outperforms in explaining the relationship between joint attributions in the input and the resulting output.

### v3

@@ A.4 Benefits to the community on improving model ability
JoPA could be used as the defense of the jailbreak attacks and help users to design better prompts to get satisfying responses. We also do the experiments to show the effectiveness of applying JoPA to detect malicious prompts on Llama-2 (7B-Chat) model, which is similar to the application done in CONTEXTCITE (Cohen-Wang et al., 2024) The results are shown below:
@@ A.6 Ablation Study
In addition, we extend the application of our framework, JoPA, to a larger model and a more challenging task, specifically Few-shot Chain-of-Thought (CoT) reasoning (Wei et al., 2022; Kojima et al., 2022), as described in Appendix A.11 and Appendix A.12. The better performance of our algorithm compared with others highlights the transferability of our framework across different model architectures and tasks.
@@ A.12 Method Generalizability on Reasoning Task
We demonstrate the generalizability of our algorithm by conducting an additional experiment on the reasoning task of Few-shot-CoT (Wei et al., 2022; Kojima et al., 2022), using the dataset AQuA (Kojima et al., 2022; Goswami et al., 2024). The following results in Figure 8 shows the remarkable performance of our method, indicating its ability to generalize well on more complex tasks like CoT.
@@ A.13 Comparison with Other Baselines
ReAGent (Zhao and Shan, 2024) addresses a different task than JoPA. It focuses on classification tasks by explaining the importance of words in inputs corresponding to the predicted label and extends this approach to generation tasks by explaining the single next predicted word. In contrast, our task focuses on explaining the relationship between the input prompt and the entire generated sentences. To compare the performance of JoPA and ReAGent on the generation task, we modify ReAGent’s explanatory target to encompass the full generation outputs rather than a single word. We conduct experiments on the Alpaca dataset using the Llama-2 (7B-Chat) model, and the results are presented in Figure 9. The experimental results demonstrate that JoPA outperforms in explaining the relationship between joint attributions in the input and the resulting output.